If angle A and angle B are linear pair, angle B + angle C = 90. Express angle A in terms of angle C.

Answer by josgarithmetic(39618)

"angle A and angle B are linear pair,"
?
A+B=180 ?

B+C=90
B=90-C
A+90-C=180
A=C+90